### v2

# Explicit bounds for the graphicality of the prime gap sequence

Abstract.
We establish explicit unconditional results on the graphic properties of the prime gap sequence. Let $p_{n}$ denote the $n$-th prime number (with $p_{0}=1$) and $\mathrm{PD}_{n}=(p_{\ell}-p_{\ell-1})_{\ell=1}^{n}$ be the sequence of the first $n$ prime gaps. Building upon the recent work by Erdős et al, which proved the graphic nature of $\mathrm{PD}_{n}$ for large $n$ unconditionally, and for all $n$ under RH, we provide the first explicit unconditional threshold such that:
  1. (1)
    For all $n\geq\exp\exp(30.5)$, $\mathrm{PD}_{n}$ is graphic.
  2. (2)
    For all $n\geq\exp\exp(34.5)$, every realization $G_{n}$ of $\mathrm{PD}_{n}$ satisfies that $(G_{n},p_{n+1}-p_{n})$ is DPG-graphic.
Our proofs utilize a more refined criterion for when a sequence is graphic, and better estimates for the first moment of large prime gaps proven through an explicit zero-free region and explicit zero-density estimate for the Riemann zeta function.
Key words and phrases:
Prime gaps, explicit estimates, graphic sequences, DPG-process.
2020 Mathematics Subject Classification:
11N05, 05C07, 11M26, 05C70.
1. Introduction
Let $p_{n}$ denote the $n$-th prime number with $p_{0}=1$. We call the sequence
$\mathrm{PD}_{n}=\left(p_{\ell}-p_{\ell-1}\right)_{\ell=1}^{n}=(1,1,2,2,4,2,\cdots,{p_{n}-p_{n-1}})$
the first $n$ prime gap sequence. A sequence of nonnegative integers $(d_{\ell})_{\ell=1}^{n}$ is said to be graphic if there exists a simple undirected graph with $n$ vertices whose degrees are exactly $d_{1},d_{2},\dots,d_{n}$. We call a simple graph on $n\geq 2$ vertices a prime gap graph if its vertex degrees are $\mathrm{PD}_{n}$.
Two key conjectures have inspired this study:
Conjecture 1.1 (Existence; Toroczkai, 2016).
For every $n\geq 2$}, there exists a prime gap graph on $n$ vertices.
Conjecture 1.2 (Structural; Toroczkai, 2016).
Every prime gap graph on $n$ vertices contains $(p_{n+1}-p_{n})/2$ independent edges.
Notably, Conjecture 1.2 implies Conjecture 1.1 (see the next subsection), linking the combinatorial properties of prime gap graphs to the distribution of primes. By merging tools from analytic number theory (to analyze prime gaps) and matching theory (to establish graph realizability and edge independence), Erdős et al. [8] proved these two conjectures hold for sufficiently large $n$ unconditionally, and for every $n$ assuming the Riemann Hypothesis (RH).
This paper establishes an explicit lower bound $n_{0}$ such that the result holds unconditionally for all $n\geq n_{0}$. Our main theorem is:
Theorem 1.3.
One may take $n_{0}=\exp\exp(30.5)$. That is, for all $n\geq\exp\exp(30.5)$, the first $n$ prime gap sequence $\mathrm{PD}_{n}$ is graphic.
While our results broadly follow the framework of [8], achieving an optimal explicit bound $n_{0}$ requires two critical refinements: a sharper criterion for graphicality of degree sequences (developed in Section 2), and a refined treatment of zero-free regions and zero-density estimates for the Riemann zeta function (detailed in Section 3).
The main bottleneck is the known explicit zero-free region for the Riemann zeta function, which impacts the explicit bounds in prime gap estimates.
1.1. DPG-process
The Degree-Preserving Network Growth (DPG) process is an algorithm to construct a sequence of graphs that iteratively extends a graph $G_{n}$ with degree sequence $D_{n}=(d_{1},\ldots,d_{n})$ to $G_{n+1}$ with $D_{n+1}=D_{n}\,\circ\,d_{n+1}:=(d_{1},\ldots,d_{n},d_{n+1})$, while preserving the degrees of the vertices of $G_{n}$. Unlike static algorithms like the Havel–Hakimi algorithm, which recompute realizations from scratch, DPG modifies $G_{n}$ through a local operation:
Degree Preserving-Step: Insert a new vertex $u$ with even degree $d_{n+1}=2\nu$.
  1. (1)
    Remove a matching of size $\nu$ from $G_{n}$ (i.e., a set of $\nu$ independent edges with no shared vertices).
  2. (2)
    Connect $u$ to all $2\nu$ endpoints of the removed edges.
Existing vertices retain their degrees, as each loses one edge from the removed matching but gains a connection to $u$.
The pair $(G_{n},d_{n+1})$ is called DPG-graphic if $G_{n}$ contains a matching of size $d_{n+1}/2$. By [9, Theorem 2.4], the sequence $D_{n}\circ d_{n+1}$ is graphic if and only if there exists a realization $G_{n}$ of $D_{n}$ such that the pair $(G_{n},d_{n+1})$ is DPG-graphic. In order to construct an infinite DPG process, one needs the requirement that every realization of $D_{n}$ contains a matching of size $d_{n+1}/2$. While this holds trivially for constant sequences (e.g., growing regular graphs), general sequences require careful analysis of matchings in $G_{n}$. To achieve this, we will rely on Vizing’s theorem and its consequence stated in Lemma 2.8.
The DPG process avoids global restructuring and ensures degree invariance, but its applicability relies on the non-trivial existence of sufficiently large matchings in intermediate graphs. It is non-trivial to construct an infinite, naturally occurring sequence $D$ whose every initial segment is graphic.
In their novel work [8], Erdős et al. demonstrated for the first time that the sequence of first $n$ prime gap graphs constitutes a non-trivial, naturally arising infinite structure (for $n$ sufficiently large and for all $n$ if RH assumed). Specifically, they established two fundamental properties: i) All initial segments of this sequence are graphic; ii) Every realization of these initial segments remains admissible for the DPG-algorithm.
Our second result makes this explicit.
Theorem 1.4.
Let $n\geq\exp\exp(34.5)$. For any realization $G_{n}$ of prime gap sequence $\mathrm{PD}_{n}$, the pair $\big(G_{n},\,p_{n+1}-p_{n}\big)$ is DPG-graphic.
1.2. Notations.
In what follows, we use the following notations. The Chebyshev function is defined by
$\psi(x)=\sum_{n\leq x}\Lambda(n),$
where $\Lambda(n)$ is the von-Mangoldt function. The notation $f=g+O^{*}(h)$ is equivalent to $|f-g|\leq h$. For $1/2\leq\sigma<1$ and $T>0$, we let $N(\sigma,T)$ be the number of non-trivial zeros $\rho=\beta+\text{i}t$ of the Riemann zeta function $\zeta(s)$ in the rectangle $\sigma<\beta<1$ and $0<t<T$.
2. Preliminaries
2.1. Some explicit estimates
Let $H_{0}\in\mathbb{R}$ be such that the RH is known to be true up to $H_{0}$. That is, all zeroes $\beta+\mathrm{i}\gamma$ of the Riemann zeta-function with $|\gamma|<H_{0}$ have $\beta=1/2$. Moreover, all of these zeroes are simple. Currently, we may take $H_{0}=3\times 10^{12}$ (see Platt and Trudgian[15, Theorem 1]). For $3/5\leq\sigma\leq 1$ and $T>H_{0}$, we have the following zero-density estimate (see Kadiri, Lumley and Ng [13, Section 5, Table 1])
$N(\sigma,T)$ $\leq{2.177}(\log T)^{5-2\sigma}T^{\frac{8}{3}(1-\sigma)}+5.663\log^{2}T$ (2.1)
$\leq{2.375}(\log T)^{5-2\sigma}T^{\frac{8}{3}(1-\sigma)}.$ (2.2)
For (2.2), we used the fact that the logarithm contribution in the first term of (2.1) is at least $\log^{3}T$ and that for $T>H_{0}$,
$\frac{5.663}{\log T}<{0.198}.$
By the zero-free region of Ford [10, Theorem 5], we have that for $|t|>3$, every non-trivial zero $\beta+\mathrm{i}t$ of the Riemann zeta function satisfies
$\beta\geq 1-\frac{c_{0}}{(\log t)^{\frac{2}{3}}(\log\log t)^{\frac{1}{3}}}:=1-\eta(t),$ (2.3)
where $c_{0}=1/57.54$. The constant $c_{0}$ has recently been improved, such as in [1, 14], with the current best result being $c_{0}=1/53.989$ by Bellotti [1, Theorem 1.2]. We therefore proceed with $c_{0}=1/53.989$ for our purpose.
Let $N(t)$ denote the number of zeros of $\zeta(s)$ in the critical strip to height $t$.
Lemma 2.1.
We have that
$N(t+1)-N(t-1)<\log t$
for all $t>10$.
Proof.
The case $t>50$ follows from Dudek [5, Lemma 2.6]. For $10<t\leq 50$, the inequality holds by direct verification of zeros of $\zeta(s)$ tabulated in the LMFDB database (https://www.lmfdb.org/zeros/zeta/). ∎
Lemma 2.2 (Delange [4]).
For any $\sigma>1$ and $t\in\mathbb{R}$ we have
$\left|\frac{\zeta^{\prime}(\sigma+\mathrm{i}t)}{\zeta(\sigma+\mathrm{i}t)}\right|<\frac{1}{\sigma-1}-\frac{1}{2\sigma^{2}}.$
Lemma 2.3 (Erdős et al. [8, Lemma 2.4]).
Let $\sigma\leq-1$ and $t\in\mathbb{R}$. Assume that either $\sigma\in 1+2\mathbb{Z}$ or $|t|\geq 1$. Then
$\left|\frac{\zeta^{\prime}(\sigma+\mathrm{i}t)}{\zeta(\sigma+\mathrm{i}t)}\right|<9+\log|\sigma+\mathrm{i}t|.$
Lemma 2.4 (Dudek [5, Lemma 2.8]).
Let $\sigma>-1$ and $t>50$. Then there exists $t_{0}\in(t-1,t+1)$ such that for every $\sigma>-1$ we have
$\left|\frac{\zeta^{\prime}(\sigma+\mathrm{i}t_{0})}{\zeta(\sigma+\mathrm{i}t_{0})}\right|<\log^{2}t+20\log t.$
Remark 1.
This lemma allows us to shift the contour slightly to a line where we have good bounds, if our contour is somewhat close to a zero.
Next we establish an explicit formula for the Chebyshev function $\psi(x)$ .
Lemma 2.5.
[…]
$\psi(x)=x-\sum_{|\operatorname{Im}\rho|<T}\frac{x^{\rho}}{\rho}+O^{*}\left(4.6\frac{x(\log x)(\log\log x)}{T}\right),$
where the sum is over the non-trivial zeros of $\zeta(s)$.
Proof.
This proof follows the framework laid out in Section 6 of [8] with some necessary modifications. We start with the truncated Perron’s formula, as presented in [8, Eq. (19)]. To avoid the potential issue that $T$ might lie close to the ordinate of a zero, we choose some $T_{0}\in(T-1,T+1)$ according to Lemma 2.4. For a non-integer $x>2$ and $c=1+1/\log x$, we have
$\psi(x)=\frac{1}{2\pi\mathrm{i}}\int_{c-\mathrm{i}T_{0}}^{{c+\mathrm{i}T_{0}}}\left(-\frac{\zeta^{\prime}(s)}{\zeta(s)}\right)\frac{x^{s}}{s}\mathrm{d}s+O^{*}\left(\sum_{n=1}^{\infty}\Lambda(n)\left(\frac{x}{n}\right)^{c}\min\left(0.501,\frac{1}{\pi T_{0}\left|\log\frac{x}{n}\right|}\right)\right).$
We estimate this error term by splitting the sum over $n$ into four parts: $S_{1}$ for $1\leq|n-x|\leq\log x$, $S_{2}$ for $\log x<|n-x|\leq x/5$, $S_{3}$ for |nx|>x/5|n-x|>x/5, and $S_{4}$ for |nx|<1|n-x|<1.
For $S_{1}$, we use the second term in the $\min$ function. By [8, Eq. (23)], we have
$S_{1}\leq 1.001\frac{x\log x}{\pi T_{0}}\sum_{\begin{subarray}{c}1\leq|x-n|\leq\log x\\ \Lambda(n)\neq 0\end{subarray}}\frac{1}{|x-n|}<0.638\frac{x\log x(\log\log x+0.6)}{T_{0}}.$
For $S_{2}$, by [8, Eqs. (24)–(26)], we have
$S_{2}\leq\frac{1}{4\log\tfrac{5}{4}}\cdot\frac{x\log x}{\pi T_{0}}\sum_{\begin{subarray}{c}\log x<|x-n|\leq x\\ \Lambda(n)\neq 0\end{subarray}}\frac{1}{|x-n|}<1.427\frac{x\log x(\log\log x-0.4)}{T_{0}}.$
For $S_{3}$, $|\log(x/n)|>\log(6/5)$ is a constant. Using Lemma 2.2, the contribution is bounded by
$S_{3}\leq\frac{1}{\pi T_{0}\log(6/5)}{\sum_{|n-x|>x/5}}\Lambda(n)\left(\frac{x}{n}\right)^{c}<\frac{ex}{\pi T_{0}\log(6/5)}\left|-\frac{\zeta^{\prime}(c)}{\zeta(c)}\right|<4.746\frac{x\log x}{T_{0}}.$
Finally for $S_{4}$, there are two $n$’s satisfying |xn|<1|x-n|<1. We use the first term in the $\min$ function, and by [8, Eq. (20)], the contribution is
$S_{4}\leq 0.501\left(\frac{x}{x-1}+\frac{x}{x}\right)\log x<1.003\log x<1.003\frac{x\log x}{T_{0}}.$
So we get
$\psi(x)=\frac{1}{2\pi\mathrm{i}}\int_{c-\mathrm{i}T_{0}}^{{c+\mathrm{i}T_{0}}}\left(-\frac{\zeta^{\prime}(s)}{\zeta(s)}\right)\frac{x^{s}}{s}\mathrm{d}s+O^{*}\left({2.065}\frac{x(\log x)(\log\log x)}{T_{0}}+{5.561}\frac{x\log x}{T_{0}}\right).$ (2.4)
The integral here is evaluated by the residue theorem, yielding
$\frac{1}{2\pi\mathrm{i}}$ $\left(\int_{c-\mathrm{i}T_{0}}^{c+\mathrm{i}T_{0}}+\int_{c+\mathrm{i}T_{0}}^{-\infty+\mathrm{i}T_{0}}+\int_{-\infty+\mathrm{i}T_{0}}^{-\infty-\mathrm{i}T_{0}}+\int_{-\infty-\mathrm{i}T_{0}}^{c-\mathrm{i}T_{0}}\right)\left(-\frac{\zeta^{\prime}(s)}{\zeta(s)}\right)\frac{x^{s}}{s}\mathrm{d}s$
$=x-\sum_{|\operatorname{Im}\rho|<T_{0}}\frac{x^{\rho}}{\rho}-\log(2\pi)-\frac{1}{2}\log(1-x^{-2}).$
By Lemma 2.3, the contribution of the third integral is
$\lim_{U\to\infty}\frac{1}{2\pi}\left|\int_{-U-\mathrm{i}T_{0}}^{-U+\mathrm{i}T_{0}}-\frac{\zeta^{\prime}(s)}{\zeta(s)}\frac{x^{s}}{s}\mathrm{d}s\right|\leq{\lim_{U\to\infty}\int_{-T_{0}}^{T_{0}}\frac{9+\log\sqrt{U^{2}+T_{0}^{2}}}{2\pi x^{U}\sqrt{U^{2}+t^{2}}}\mathrm{d}t}=0.$
By Lemma 2.3 and Lemma 2.4, the contribution of the second integral is
$\frac{1}{2\pi}\left|\int_{c+\mathrm{i}T_{0}}^{-\infty+\mathrm{i}T_{0}}\left(-\frac{\zeta^{\prime}(s)}{\zeta(s)}\right)\frac{x^{s}}{s}\mathrm{d}s\right|\leq\frac{{\log^{2}T+20\log T}}{2\pi T_{0}}\int_{-\infty}^{c}x^{\sigma}\mathrm{d}\sigma\leq\frac{ex(\log x+20)}{2\pi T_{0}}.$
The fourth integral obeys the same bound. Plugging all these into (2.4) we have
$\psi(x)=x$ $-\sum_{|\operatorname{Im}\rho|<T_{0}}\frac{x^{\rho}}{\rho}-\log(2\pi)-\frac{1}{2}\log(1-x^{-2})+O^{*}\left(\frac{ex(\log x+20)}{\pi T_{0}}\right)$ (2.5)
$+O^{*}\left(2.065\frac{x(\log x)(\log\log x)}{T_{0}}+{5.561}\frac{x\log x}{T_{0}}\right).$
By Lemma 2.1, we adjust the $\rho$-sum to $|\operatorname{Im}\rho|<T$ at the cost of an error of
$\left|\sum_{|\operatorname{Im}\rho-T|\leq 1}\frac{x^{\rho}}{\rho}\right|\leq\sum_{|\operatorname{Im}\rho-T|\leq 1}\frac{x}{T-1}\leq\frac{2x\log T}{T-1}.$ (2.6)
Note that for $x>10^{18},\,T>50$, we have
$1<0.269\log\log x,\quad 1<0.007(\log x)(\log\log x),\quad\text{and}\ \frac{1}{T-1}\leq 1.021\frac{1}{T}.$
Combining (2.5) and (2.6) we have, for a non-integer $x$,
$\psi(x)=x-\sum_{|\operatorname{Im}\rho|<T}\frac{x^{\rho}}{\rho}+O^{*}\left(4.6\frac{x(\log x)(\log\log x)}{T}\right).$
This completes the proof of this lemma. ∎
Remark 2.
Note that Lemma 2.5 gives an explicit version of a result due to Goldston [11, Eq. (4)]; see also another explicit version in Cully-Hugill [2, Theorem 2] for half odd integers.
Let $\pi(x)$ be the prime-counting function defined to be the number of primes less than or equal to $x$.
Lemma 2.6 (Dusart[6]).
We have
$\frac{x}{\log x}\left(1+\frac{1}{\log x}\right)\leq\pi(x)\leq\frac{x}{\log x}\left(1+\frac{1}{\log x}+\frac{2.51}{(\log x)^{2}}\right).$
The first inequality holds for all $x\geq 599$ and the second one for $x\geq 355991$.
2.2. Criteria for sequences being graphic
To address this problem, we draw upon some knowledge from graph theory. Many results provide necessary and sufficient conditions for a nonnegative sequence to be graphic. One classic and well-known theorem is the Erdős–Gallai criterion [7]: Given $\left(d_{i}\right)_{i=1}^{n}\in\mathbb{N}^{n}$ such that $d_{1}\geq d_{2}\geq\cdots\geq d_{n}\geq 0$ and the sum of all terms is even, then there exists a (simple undirected) graph with $n$ vertices and degrees $d_{1},\dots,d_{n}$ if and only if
$\sum_{i=1}^{k}d_{i}\leq k(k-1)+\sum_{i=k+1}^{n}\min(k,d_{i})$
for all $k\in\{1,\dots,{n-1}\}$.
In fact, the results of Zverovich and Zverovich [20], and later Tripathi and Vijay [18], indicate that it is unnecessary to verify the above inequality for all values of $1\leq k\leq n$.
Theorem 2.7.
Let $\left(d_{i}\right)_{i=1}^{n}\in\mathbb{N}^{n}$ be a nonincreasing sequence with $d_{1}\geq d_{2}\geq\cdots\geq d_{n}\geq 0$. Then $\left(d_{i}\right)_{i=1}^{n}$ is graphic if and only if $\sum_{i=1}^{n}d_{i}$ is even and
$\sum_{i=1}^{k}d_{i}\leq k(k-1)+\sum_{i=k+1}^{n}\min(k,d_{i})$
holds for $k=m$ and for all $k<m$ such that $d_{k}>d_{k+1}$, where $m=\max\{i:d_{i}\geq i\}$.
This result significantly reduces the number of $k$-values that need to be checked in the original Erdős–Gallai criterion, especially when $n$ is sufficiently large. This reduction plays an important role in our setting.
2.3. Application of Vizing’s theorem
A well-known theorem by Vizing [19] states that a simple graph with maximal degree $\Delta$ admits a proper edge coloring with $\Delta+1$ colors. This leads to the following lemma, which establishes a condition for the existence of a large matching.
Lemma 2.8 (Erdős et al.[8, Lemma 2.1]).
Let $G$ be a simple graph on $n$ vertices with degrees $d_{1},\dots,d_{n}$. Let $\delta\geq 1$ be an integer, and let $d\geq 2$ be an even integer satisfying
$\delta d\leq\sum_{d_{i}<\delta}d_{i}-\sum_{d_{i}\geq\delta}d_{i}.$
Then $G$ has a matching of size $d/2$.
3. First moment of large prime gaps
Let $N=4\delta x$ throughout this section. For simplicity, we define the integral
$\mathcal{I}=\mathcal{I}(x,T):=\int_{x}^{2x}\left|\sum_{|\operatorname{Im}\rho|<T}y^{\rho}C(\rho)\right|^{2}\mathrm{d}y$
where
$C(\rho)=\frac{1-(1+\delta)^{\rho}}{\rho}$
with $0<\delta\leq 1$. Following essentially the method of Heath-Brown [12], we will bound the first moment of large prime gaps by the integral $\mathcal{I}$.
Lemma 3.1.
Let $x\geq\exp{10^{3}}$ and $N>8\log^{2}x.$ We have
$\sum_{\begin{subarray}{c}x\leq p_{\ell}\leq 2x\\ p_{\ell+1}-p_{\ell}>N\end{subarray}}\left(p_{\ell+1}-p_{\ell}\right)\leq\max_{\begin{subarray}{c}x\leq p_{\ell}\leq 2x\end{subarray}}\left(p_{\ell+1}-p_{\ell}\right)+8.1\frac{1}{\delta^{2}x^{2}}\mathcal{I}(x,2\delta^{-1}\log^{2}x).$
Proof.
Assume that the prime $p_{\ell}\in[x,2x]$ satisfies ${p_{\ell+1}}-p_{\ell}\geqslant N.$ There is at most one $p_{\ell}$ such that $(p_{\ell}+p_{\ell+1})/2>2x$, so we consider its contribution separately and assume that $(p_{\ell}+p_{\ell+1})/2\leqslant 2x$. Then, for any
$y\in(p_{\ell},(p_{\ell}+p_{\ell+1})/2)\subset(x,2x),$
the interval
$y,y+\delta y]\subset[y,y+N/2]\subset(p_{\ell},p_{\ell+1})$
is free of primes. The difference $\psi(y+\delta y)-\psi(y)$ therefore only includes contribution from higher prime powers $p^{k}$ ($k\geq 2$), which gives
$\sum_{\begin{subarray}{c}p^{k}\in(y,y+\delta y]\\ k\geq 2\end{subarray}}\log p$ $\leq\sum_{2\leq k\leq\frac{\log(y+\delta y)}{\log 2}}\,\sum_{\begin{subarray}{c}y^{1/k}<p\leq(y+\delta y)^{1/k}\end{subarray}}\log p$
$<\sum_{k\geq 2}\frac{1}{k}\log(y+\delta y)\left(\pi\big((y+\delta y)^{1/k}\big)-\pi\big(y^{1/k}\big)\right)$
$<\log(y+\delta y)\sum_{k\geq 2}\frac{1}{k}y^{1/k}\left((1+\delta)^{1/k}-1\right)$
[…]
for all $y\geq x\geq 1.6\times 10^{8}$. Now we have $|\psi(y+\delta y)-\psi(y)-\delta y|>0.999\delta y$ holds on $y\in(p_{\ell},(p_{\ell}+p_{\ell+1})/2).$ Squaring and integrating, we obtain
$\int_{p_{\ell}}^{(p_{\ell}+p_{\ell+1})/2}|\psi(y+\delta y)-\psi(y)-\delta y|^{2}\mathrm{d}y>0.499\delta^{2}x^{2}(p_{\ell+1}-p_{\ell}).$
Note that the set of points where $y+\delta y$ or $y$ takes an integer value is countable and thus of Lebesgue measure zero, making its contribution to the above integral zero. It follows from Lemma 2.5 that
$|\psi(y+\delta y)-\psi(y)-\delta y|$ $=\left|\sum_{|\operatorname{Im}\rho|<T}\frac{y^{\rho}-(y+\delta y)^{\rho}}{\rho}+E(y+\delta y,T)-E(y,T)\right|$
$\leq\left|\sum_{|\operatorname{Im}\rho|<T}y^{\rho}C(\rho)\right|+E(T,x).$
With the choice of $T=2\delta^{-1}\log^{2}x<x$, the error term $E(T,x)$ satisfies
$E(T,x)$ $\leq{4.6}\frac{4x\log(4x)\log\log(4x)+2x\log(2x)\log\log(2x)}{T}$
$\leq\frac{x}{T}\log^{2}x=\frac{1}{2}\delta x,$
for all $x\geq 3.9\times 10^{59}$. On using the simple inequality $(A+B)^{2}\leq 2A^{2}+2B^{2}$, we have
$0.499\delta^{2}x^{2}(p_{\ell+1}-p_{\ell})<2\int_{p_{\ell}}^{(p_{\ell}+p_{\ell+1})/2}\left|\sum_{|\operatorname{Im}\rho|<T}y^{\rho}C(\rho)\right|^{2}\mathrm{d}y+\frac{1}{4}\delta^{2}x^{2}(p_{\ell+1}-p_{\ell}).$
Summing over all such primes $p_{\ell}$, and bounding the possible single exceptional $p_{\ell}$ by taking maximum, we obtain the desired result. ∎
Remark 3.
Lemma 3.1 is also established in [8, Section 5]. Here, we present a reformulation and a minor modification to it. Specifically, the conclusion of Lemma 2.1 ensures validity for $T<x$, which allows us to take a smaller imaginary part.
The main task now is to give the non-trivial estimate of Integral $\mathcal{I}$ explicitly. This can be done by following the same idea as in Saffari and Vaughan [17, Lemma 5].
Proposition 3.2 (Bounds for Integral $\mathcal{I}$).
Suppose that there are positive constants $c_{1},A$ (with $2\leq A\leq 4$}) such that
$N(\sigma,t)\leq c_{1}t^{A(1-\sigma)}(\log t)^{5-2\sigma},\quad\left(t\geq 2,\,\,\sigma\geq 1-1/A\right).$ (3.1)
Suppose further that the Riemann zeta function has no zeros $\rho=\beta+\mathrm{i}\gamma$ in the region
$\beta\geq 1-\eta({\gamma}),\quad{\gamma}\geq 2,$ (3.2)
with $\eta({\gamma})<1/A$. Then, whenever $x\geq\exp{10^{3}}$, $0<\alpha<2/A$ and $10<T\leq x^{2/A-\alpha}/\log x$, we have
$\mathcal{I}(x,T)\leq C\,\delta^{2}x^{3-A\alpha\eta(T)}\log^{4}x,$
where the constant
$C=C(A,\alpha,c_{1})=4\times 1836\,\frac{c_{1}}{A\alpha}\left(\frac{2}{A}-\alpha\right)^{4}.$
Proof.
We can first bound this integral as follows
$\begin{split}\mathcal{I}&\leq\int_{1}^{2}\int_{\frac{vx}{2}}^{2vx}\left|\sum_{|\operatorname{Im}\rho|\leq T}y^{\rho}C(\rho)\right|^{2}\mathrm{d}y\,\mathrm{d}v\\ &\leq\int_{1}^{2}\sum_{|\operatorname{Im}\rho|\leq T}\sum_{|\operatorname{Im}\rho^{\prime}|\leq T}C(\rho)\overline{C(\rho^{\prime})}\frac{2^{1+\rho+\overline{\rho^{\prime}}}-2^{-1-\rho-\overline{\rho^{\prime}}}}{1+\rho+\overline{\rho^{\prime}}}x^{1+\rho+\overline{\rho^{\prime}}}v^{1+\rho+\overline{\rho^{\prime}}}\mathrm{d}v\\ &\leq x\sum_{|\operatorname{Im}\rho|\leq T}\sum_{|\operatorname{Im}\rho^{\prime}|\leq T}\left|x^{\rho+\overline{\rho^{\prime}}}C(\rho)\overline{C(\rho^{\prime})}\right|\cdot\left|\frac{2^{1+\rho+\overline{\rho^{\prime}}}-2^{-1-\rho-\overline{\rho^{\prime}}}}{1+\rho+\overline{\rho^{\prime}}}\right|\cdot\left|\frac{2^{2+\rho+\overline{\rho^{\prime}}}-1}{2+\rho+\overline{\rho^{\prime}}}\right|.\end{split}$
By writing $\rho=\beta+\mathrm{i}\gamma,\rho^{\prime}=\beta^{\prime}+\mathrm{i}\gamma^{\prime}$, and using the elementary inequality $2|z_{1}z_{2}|\leq|z_{1}|^{2}+|z_{2}|^{2}$, the above is
$\leq\frac{1}{2}x\sum_{\begin{subarray}{c}\rho=\beta+\mathrm{i}\gamma\\ |\gamma|\leq T\end{subarray}}\sum_{\begin{subarray}{c}\rho^{\prime}=\beta^{\prime}+\mathrm{i}\gamma^{\prime}\\ |\gamma^{\prime}|\leq T\end{subarray}}\left(x^{2\beta}|C(\rho)|^{2}+x^{2\beta^{\prime}}|C(\rho^{\prime})|^{2}\right)\left(\frac{(2^{3}+1)(2^{4}+1)}{1+|\gamma-\gamma^{\prime}|^{2}}\right),$
which gives
$\mathcal{I}$ $\leq 153x\sum_{\begin{subarray}{c}\rho=\beta+\mathrm{i}\gamma\\ |\gamma|\leq T\end{subarray}}x^{2\beta}|C(\rho)|^{2}\sum_{\begin{subarray}{c}\rho^{\prime}=\beta^{\prime}+\mathrm{i}\gamma^{\prime}\\ |\gamma^{\prime}|\leq T\end{subarray}}\frac{1}{1+|\gamma-\gamma^{\prime}|^{2}}$
$\leq 4\times 153x\sum_{\begin{subarray}{c}\rho=\beta+\mathrm{i}\gamma\\ \beta\geq\frac{1}{2},10<\gamma\leq T\end{subarray}}x^{2\beta}|C(\rho)|^{2}\sum_{\begin{subarray}{c}\rho^{\prime}=\beta^{\prime}+\mathrm{i}\gamma^{\prime}\\ |\gamma^{\prime}|\leq T\end{subarray}}\frac{1}{1+|\gamma-\gamma^{\prime}|^{2}}$
by the symmetry of the double summation and the symmetric distribution of zeros. We divide the summation over $\gamma^{\prime}$ by their distance from $\gamma$, and apply Lemma 2.1 to get that the inner sum of above is
$\leq\sum_{k\geq 0}\,\sum_{k\leq|\gamma^{\prime}-\gamma|\leq k+1}\frac{1}{1+|\gamma^{\prime}-\gamma|^{2}}\leq\sum_{k\geq 0}\frac{1}{1+k^{2}}\sum_{k\leq|\gamma^{\prime}-\gamma|\leq k+1}1$
$\leq\sum_{k\geq 0}\frac{\log(\gamma+k)}{1+k^{2}}\leq\sum_{k=0}^{\lfloor\gamma\rfloor}\frac{\log(2\gamma)}{1+k^{2}}+\sum_{k>\gamma}\frac{\log(2k)}{k^{2}}\leq\left(\frac{1}{2}+\frac{\pi}{2}\coth(\pi)\right)\log(2\gamma)+\frac{\log(2\gamma)+1}{\gamma}$
$\leq 3\log\gamma$
for all $\gamma>10$. Using the integral representation
$C(\rho)=\int_{1}^{1+\delta}x^{\rho-1}\mathrm{d}x,$
we apply the triangle inequality for complex-valued integrals to get
$|C(\rho)|\leq\int_{1}^{1+\delta}|x^{\rho-1}|\mathrm{d}x=\int_{1}^{1+\delta}x^{\beta-1}\mathrm{d}x\leq\int_{1}^{1+\delta}1\,\mathrm{d}x=\delta.$
Thus we have
$\begin{split}&\mathcal{I}\leq 1836x\delta^{2}\log T\sum_{\begin{subarray}{c}\rho=\beta+\mathrm{i}\gamma\\ \beta\geq\frac{1}{2},10<\gamma\leq T\end{subarray}}x^{2\beta}.\end{split}$ (3.3)
Let
$S=\sum_{\begin{subarray}{c}\rho=\beta+\mathrm{i}\gamma\\ \beta\geq\frac{1}{2},10<\gamma\leq T\end{subarray}}x^{2\beta}=S_{1}+S_{2},$
with
$S_{1}:=\sum_{\begin{subarray}{c}\rho=\beta+\mathrm{i}\gamma\\ \beta\geq\frac{1}{2},10<\gamma\leq T\end{subarray}}\left(x^{2\beta}-x\right),\quad\text{and}\quad S_{2}:=x\sum_{\begin{subarray}{c}\rho=\beta+\mathrm{i}\gamma\\ \beta\geq\frac{1}{2},10<\gamma\leq T\end{subarray}}1.$
By Lemma 2.1, we have
$S_{2}$ $\leq xN\left(T\right)\leq xT\log T.$ (3.4)
By the zero-free region (3.2), we have
$S_{1}$ $=\sum_{\beta\geq\frac{1}{2},\,10<\gamma\leq T}\int_{1/2}^{\beta}2x^{2\sigma}\log x\ \mathrm{d}\sigma$
$\leq\int_{1/2}^{1-\eta(T)}2x^{2\sigma}(\log x)\,N(\sigma,T)\,\mathrm{d}\sigma.$
We split the integral above into two parts, one over the interval $1/2\leq\sigma\leq 1-1/A$, and another over $1-1/A\leq\sigma\leq 1-\eta(T)$. By Lemma 2.1, the contribution of first interval is
$\leq T(\log T)(x^{2-\frac{2}{A}}-x).$ (3.5)
And the contribution over $1-1/A,1-\eta(T)]$ is estimated by the zero-density estimate (3.1), which gives
$\leq 2c_{1}(\log x)\int_{1-1/A}^{1-\eta(T)}x^{2\sigma}T^{A(1-\sigma)}(\log T)^{5-2\sigma}\,\mathrm{d}\sigma$
$\leq 2c_{1}x^{2}(\log x)(\log^{3}T)\,\frac{K^{-\eta(T)}}{\log K},$ (3.6)
where
$K=\frac{x^{2}}{T^{A}\log^{2}T}.$
In the end, since it is assumed that $K\geq x^{A\alpha}>1$ and $\eta(T)<1/A$, by combining (3.4), (3.5) and (3) we obtain
$S\cdot(\log T)$ $\leq\left(\frac{2}{A}-\alpha\right)^{2}x^{2-\alpha}\log^{2}x+\frac{2c_{1}}{A\alpha}\left(\frac{2}{A}-\alpha\right)^{4}x^{2-A\alpha\eta(T)}\log^{4}x.$ (3.7)
By inserting (3.7) into (3.3), we complete the proof.
With our choice $T=2\delta^{-1}\log^{2}x$, we have
$\mathcal{I}(x,2\delta^{-1}\log^{2}x)\leq C\delta^{2}x^{3-A\alpha\eta(2\delta^{-1}\log^{2}x)}\log^{4}x.$
This is valid for $N\geq 8x^{1+\alpha-2/A}\log^{3}x.$ Applying the best known explicit results on the zero-density estimate (2.2) and the zero-free region (2.3), we find that $A=8/3,\,c_{1}={2.375}$, and
$\eta(T)=\eta\left(\frac{8x\log^{2}x}{N}\right)=\frac{1}{53.989}\left(\log\frac{8x\log^{2}x}{N}\right)^{-2/3}\left(\log\log\frac{8x\log^{2}x}{N}\right)^{-1/3}.$
Proposition 3.2, together with Lemma 3.1 then gives
$\sum_{\begin{subarray}{c}x\leq p_{\ell}\leq 2x\\ p_{\ell+1}-p_{\ell}>N\end{subarray}}\left(p_{\ell+1}-p_{\ell}\right)$ $\leq\max_{\begin{subarray}{c}x\leq p_{\ell}\leq 2x\end{subarray}}\left(p_{\ell+1}-p_{\ell}\right)+8.1Cx^{1-\frac{8}{3}\alpha\eta(8N^{-1}x\log^{2}x)}\log^{4}x.$ (3.8)
Hence, we establish the following theorem, which will be primarily used to prove Theorem 1.3 and Theorem 1.4.
Theorem 3.3.
Let $x\geq\exp(4\times 10^{3}),\,N\geq 8x^{1/4+\alpha}\log^{3}x$ with $0<\alpha<3/4$. Then
$S_{N}(x):=\sum_{\begin{subarray}{c}p_{\ell}\leq x\\ p_{\ell+1}-p_{\ell}>N\end{subarray}}\left(p_{\ell+1}-p_{\ell}\right)\leq M(x)\log x+8.7C\,x^{1-\frac{8}{3}\alpha\eta\left(4N^{-1}x\log^{2}x\right)}\log^{4}x,$
where $M(x)$ is the largest prime gaps up to $x$.
Remark 4.
This theorem can be compared with Theorem 1.8 in [8], where the latter assumes RH. Informally, our result serves as its unconditional counterpart, albeit with a weaker exponent in the upper bound. This compromise requires that the subsequent main Theorem 1.3 is vaild only for sufficiently large $n$.
Proof.
First note that
$\sum_{\begin{subarray}{c}y/2<p_{\ell}\leq y\\ p_{\ell+1}-p_{\ell}>N\end{subarray}}(p_{\ell+1}-p_{\ell})=0,$
if $N\geq y$. Let $k$ be the integer such that $2^{-k}x<N\leq 2^{-k+1}x$, then we have $S_{N}(2^{-k}x)=0$ and
$2^{-k}x\geq 2^{-1}N\geq 4x^{\frac{1}{4}+\alpha}\log^{3}x\geq\exp{10^{3}}.$
Let $\eta(t)$ be as in (2.3) for $t>3$. By the dyadic argument and (3.8) we have
$S_{N}(x)$ $=\sum_{\begin{subarray}{c}\frac{x}{2}\leq p_{\ell}\leq x\\ p_{\ell+1}-p_{\ell}>N\end{subarray}}(p_{\ell+1}-p_{\ell})+\cdots+\sum_{\begin{subarray}{c}\frac{x}{2^{k}}\leq p_{\ell}\leq\frac{x}{2^{k-1}}\\ p_{\ell+1}-p_{\ell}>N\end{subarray}}(p_{\ell+1}-p_{\ell})$
$\leq M(x)\log x+8.1C\sum_{j=1}^{k}\left(\log\frac{x}{2^{j}}\right)^{4}\left(\frac{x}{2^{j}}\right)^{1-\frac{8}{3}\alpha\eta\left(8N^{-1}\frac{x}{2^{j}}\log^{2}(\frac{x}{2^{j}})\right)}.$ (3.9)
Since $\eta(t)$ is decreasing, we obtain
$\eta\left(\frac{8x\log^{2}(\frac{x}{2^{j}})}{2^{j}N}\right)\geq\eta\left(\frac{4x\log^{2}x}{N}\right),$
and with $2^{-k}x\geq N/2$ to get
$\eta\left(\frac{8x\log^{2}(\frac{x}{2^{j}})}{2^{j}N}\right)$ $\leq\eta\left(\frac{8x\log^{2}(\frac{x}{2^{k}})}{2^{k}N}\right)\leq\eta\left(4\log^{2}\frac{N}{2}\right)\leq\eta(4)<\frac{1}{40}.$
Then the summation on the right-hand side of (3) is
$\leq x^{1-\frac{8}{3}\alpha\eta\left(4N^{-1}x\log^{2}x\right)}\log^{4}x\,\sum_{j\geq 1}\left(\frac{1}{2^{j}}\right)^{1-\frac{8}{3}\times\frac{3}{4}\times\frac{1}{40}}$
$\leq 1.074\,x^{1-\frac{8}{3}\alpha\eta\left(4N^{-1}x\log^{2}x\right)}\log^{4}x,$
where the infinite sum over $j$ is convergent to $1.0731\dots$. This completes the proof. ∎
4. Proof of Theorem 1.3
Thanks to Theorem 2.7, to prove Theorem 1.3, it suffices to check the Erdős–Gallai inequalities where the sequence jumps. We define
$k_{N}(x):=|\{\ell\in\mathbb{N}:\,p_{\ell}\leq x,\,p_{\ell+1}-p_{\ell}\geq N\}|,$
and write $k_{N}=k_{N}(p_{n})$ for simplicity.
Thus we are required to establish
$\sum_{\begin{subarray}{c}1\leq\ell\leq n\\ p_{\ell+1}-p_{\ell}>N\end{subarray}}(p_{\ell+1}-p_{\ell})\leq k_{N}(k_{N}-1)+n-k_{N}$
for all $N\geq 2$, which is equivalent to establishing
$S_{N}(p_{n})\leq(k_{N}-1)^{2}+n-1$
for all $1\leq k_{N}\leq n$.
4.1. Case 1: $k_{N}\geq\sqrt{p_{n}}+1$
This is the trivial case, since
$S_{N}(p_{n})\leq p_{n}\leq(k_{N}-1)^{2}.$
4.2. Case 2: $k_{N}\leq\sqrt{p_{n}}$
In this case, we choose $\alpha=0.249$, and let
$N^{\prime}=\frac{p_{n}^{1/2}}{3\log p_{n}}.$
It is easy to see that $N^{\prime}\geq 8p_{n}^{1/4+\alpha}\log^{3}p_{n}$ for $n\geq\exp\exp(10.8)$. Then we have by Theorem 3.3 that
$S_{N}(p_{n})$ $=\sum_{\begin{subarray}{c}1\leq\ell\leq n\\ N\leq p_{\ell+1}-p_{\ell}<N^{\prime}\end{subarray}}(p_{\ell+1}-p_{\ell})+S_{N^{\prime}}(p_{n})$
$\leq N^{\prime}k_{N}+M(p_{n})\log p_{n}+8.7C\,p_{n}^{1-\frac{8}{3}\alpha\eta\left(\frac{4p_{n}\log^{2}p_{n}}{N^{\prime}}\right)}\log^{4}p_{n}.$ (4.1)
By the lower bound on $\pi(p_{n})$ (Lemma 2.6), we have
$\frac{p_{n}}{\log p_{n}}<\frac{p_{n}}{\log p_{n}}\big(1+\frac{1}{\log p_{n}}\big)-1\leq n-1.$
So the first term in (4.1) is
$N^{\prime}k_{N}\leq\frac{p_{n}}{3\log p_{n}}<\frac{n-1}{3}.$ (4.2)
Using the fact that there is at least one prime between $n^{90}$ and $(n+1)^{90}$ for all $n\geq 1$ (see Cully-Hugill and Johnston [3, Theorem 1.4]), we crudely bound
$M(p_{n})\leq 180\,\left(p_{n}^{1/90}+2\right)^{89}\leq\frac{p_{n}}{3\log^{2}p_{n}}$}
for $n\geq\exp\exp(10)$. Thus the second term in (4.1) is
$M(p_{n})\log p_{n}\leq\frac{p_{n}}{3\log^{2}p_{n}}\log p_{n}=\frac{p_{n}}{3\log p_{n}}<\frac{n-1}{3}.$ (4.3)
For the third term in (4.1), we write $t=\log\log p_{n}$. Numerical computations show that the following results
$\log(3\times 8.7C)+5t<\frac{8\times 0.249}{3\times 53.989}\cdot\frac{e^{t}}{L_{1}^{2/3}\log^{1/3}L_{1}}$
hold for all $t\geq 30.5$, where
$L_{1}=L_{1}(t)=\frac{1}{2}e^{t}+3t+\log{12},$
and
$C=4\times{1836}\,\frac{c_{1}}{A\alpha}\left(\frac{2}{A}-\alpha\right)^{4}{\approx 1654.9\leq 1655},$
with $\alpha=0.249$, $A=8/3$, $c_{1}={2.375}$ and $c_{0}=1/53.989$. This implies that
$8.7C\,p_{n}^{1-\frac{8}{3}\alpha\eta\left(\frac{4p_{n}\log^{2}p_{n}}{N^{\prime}}\right)}\log^{4}p_{n}<\frac{p_{n}}{3\log p_{n}}<\frac{n-1}{3}$ (4.4)
for $n\geq\exp\exp({30.5})$. Therefore by (4.2)–(4.4), it is immediate that
$S_{N}(p_{n})<3\cdot\frac{n-1}{3}=n-1.$
Combining these two cases, we have thus completed the proof of Theorem 1.3.
5. About the DPG-process
We have proved that the sequence of first $n$ prime gaps (with $p_{0}=1$) is graphic for $n\geq\exp\exp(30.49)$. Let $G_{n}$ be the realization. By Erdős et al.[9, Theorem 2.5], to prove that $(p_{\ell}-p_{\ell-1})_{\ell=1}^{n}$ is DPG-graphic, it suffices to show that $G_{n}$ has $(p_{n+1}-p_{n})/2$ independent edges. Applying Lemma 2.8, we are required to exhibit an integer $N\geq 2$ satisfying
$N(p_{n+1}-p_{n})+2\sum_{\begin{subarray}{c}1\leq\ell\leq n\\ p_{\ell}-p_{\ell-1}>N\end{subarray}}\left(p_{\ell}-p_{\ell-1}\right)<p_{n},$
for all $n\geq n_{0}$ (This is [8, Eq. (11)]). In [5], the author shows that for $x>\exp\exp(33.3)$ one has the bound $p_{n+1}-p_{n}<3x^{2/3}+3x^{1/3}+2$ when $x<p_{n}<2x$. Choose $\alpha=1/13$ such that
$N=\frac{1}{12}p_{n}^{1/3}\geq 8p_{n}^{1/4+\alpha}\log^{3}p_{n}$
for $n\geq\exp\exp(8.5)$, then we have
$N(p_{n+1}-p_{n})\leq\frac{3.1}{12}p_{n}^{1/3+2/3}<\frac{1}{3}p_{n}.$
By Theorem 3.3, we also need the inequality
$M(p_{n})\log p_{n}+8.7Cp_{n}^{1-\frac{8}{3}\alpha\eta(48p_{n}^{2/3}\log^{2}p_{n})}\log^{4}p_{n}<\frac{1}{3}p_{n}.$
By (4.3), it suffices to show that
$8.7Cp_{n}^{1-\frac{8}{3}\alpha\eta(48p_{n}^{2/3}\log^{2}p_{n})}\log^{4}p_{n}<\frac{1}{4}p_{n}<\frac{1}{3}p_{n}(1-\frac{1}{\log p_{n}}),$
for all $n\geq\exp\exp(10)$. This can be reduced to
$\log(4\times 8.7C)+4t<\frac{8}{3\times 13\times 53.989}\times\frac{e^{t}}{L_{2}^{2/3}\log^{1/3}L_{2}},$
where
$L_{2}=L_{2}(t)=\frac{2}{3}e^{t}+2t+\log 48,$
and
$C=4\times{1836}\,\frac{c_{1}}{A\alpha}\left(\frac{2}{A}-\alpha\right)^{4}{\approx 17451.4\leq 17452},$
by writing $t=\log\log p_{n}$. By numerical verification, this is true for all $t\geq 34.5$. Therefore, we conclude the proof of Theorem 1.4.
Acknowledgements. The authors would like to thank Prof. Gergely Harcos for suggesting the problem and his valuable suggestions on this topic. We also thank Daniel Johnston for pointing out the references [2, 3]. K. Aggarwal thanks the Rényi Institute for funding and support where most of the work was done. R. Frot also thanks the Rényi and the Erdős Center for funding and support, which were provided in part by the Rényi Intézet Lendület Automorphic Research Group. H. Gou and H. Wang are grateful to Prof. Jianya Liu for his help and encouragement. H. Wang also gratefully acknowledges the Rényi Institute for its excellent research environment, which enabled the preparation of a portion of this work, as well as the China Scholarship Council for supporting her studies there.